P4 Sunday, July 19, 2020 and (1-P) 6:03 PM probabilities Р ML respectively (20%) Suppose that a signals that takes on valuesFind the Maximum likelihood estimation, maximum a posteriori estimation, and the mean absolute error. Provided is the mean squared error.
Answer #1

ANSWER:

Given RIN(8, 2) with paffils) PL3=1) =P = P(s =4)=1-P Posterior distribution of the signal sent ; is = f(ris = 1) PCS=1) f(18P-ci-pler P +(1-P) er. .: Posterior Mean P. (i-pler 4 cr-per